OpenStudy (anonymous):

Find f(-7) if f(x) = 2x + 8. -6 2 3 22 Can you explain again how to work this out im having troube @ilfy214

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

\[f(x) = 2x + 8\]Start with your equation and then plug in the "x" \[x = -7\]\[f(-7) = 2(-7) + 8\]So now you multiply and add! What do you get?

OpenStudy (anonymous):

does x equal -7??

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Now finish your equation...

OpenStudy (anonymous):

f(-7)=-6 i think

OpenStudy (anonymous):

YES! Correct \[f(-7) = 2(-7) + 8\]\[f(-7) = -14 + 8\]\[f(-7) = -6\]

OpenStudy (anonymous):

is tht the awnser @ilfy214 sry i was gone for a bit

OpenStudy (anonymous):

@ilfy214

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

kk tyty

OpenStudy (anonymous):

Anytime!